How many of the possible 5-card hands from a standard 52-card deck would consist of the following cards?
a) 4 clubs and 1 non-club
b) 2 face cards and 3 non-face cards
c) 2 red cards, 2 clubs and a spade

Answer by stanbon(57412)

How many of the possible 5-card hands from a standard 52-card deck would consist of the following cards?
a) 4 clubs and 1 non-club
Pick 4 clubs: 13C4 = 715 ways
Pick a non-club: 39
# of possible hands: 715*39 = 27885
-----------------------------------------
b) 2 face cards and 3 non-face cards
Pick 2 face cards: 66
Pick 3 non-face cards: 40C3 = 9880
# of possible hands: = 66*9880 = 652080
-----------------------------------------
c) 2 red cards, 2 clubs and a spade
Pick 2 red card: 26C2 = 325
Pick 2 clubs: 13C2 = 78
Pick a spade: 13
# of possible hands: 325*78*13 = 329550
=========================================
Cheers,
Stan H.